OpenStudy (anonymous):

A 2.0-gram bullet is shot into a tree stump. It enters at a speed of 3.00 x 104 cm/s and comes to rest after having penetrated 0.05 m in to the stump. What was the average force during the impact? Show all calculations leading to an answer.

OpenStudy (theeric):

What does Newton's second law tell us?

OpenStudy (anonymous):

i think its what ever is in motion stays in motion

OpenStudy (theeric):

That's part of the first law!

OpenStudy (anonymous):

opps lol i get my laws mixed up

OpenStudy (theeric):

Okay! So, do a quick Google search! If half a brain isn't up to the task, going for a short walk can do wonders.

OpenStudy (anonymous):

Newton's second law of motion pertains to the behavior of objects for which all existing forces are not balanced. The second law states that the acceleration of an object is dependent upon two variables - the net force acting upon the object and the mass of the object.

OpenStudy (theeric):

Alright! Thank http://www.physicsclassroom.com/class/newtlaws/Lesson-3/Newton-s-Second-Law . The important part is just this: \(F_{net}=ma\). So, the averages are proportional by inertial mass as well. \(\bar F = m\bar a\) \(v_f^2=v_i^2+\frac12\bar ad\Rightarrow \bar a=\dfrac{-2v_i^2}{d}\) So, \(\bar F = -\dfrac{2m\ v_i^2}{d}\).

OpenStudy (anonymous):

so how dod you get thatim very curios haha

OpenStudy (dan815):